# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ I.1 Bounds on Number of Floppy Modes
Figure 5: Average number of floppy modes in randomly generated lattices. a, Results for different system sizes normalized as $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, such that the lower bound will coincide. b, Normalized umber of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
To maximize the number $F$ of floppy mode, we minimize the number $R$ of rigid chains and maximize the number $L$ of loops. The minimal value of $R$ is zero. We now derive two upper bounds on $L$. The first stems from the fact that each loop surrounds at least one internal corner node, thus $L\leq\frac{N+2-P}{2}$, which leads to $F\leq N_{1}+1$.
@@
For a given lattice of size and shape, the system becomes more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with $E-1$ bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in the expression for the number of modes, we obtain the lower bound given in Eq. (1).
@@
I.2 Multi-Layer Matrix-Vector Multiplication
We have demonstrated the matrix-vector multiplication in a single-layer metamaterial. Here, we further show a larger matrix-vector multiplication in a three-layer metamaterial of four inputs in the middle layer and 12 outputs in the side layers. These frustrated loops similarly relate the input and output vectors. We experimentally apply a single input and measure the transmission matrix $A$ from the slopes of the input-output curves (Fig. 6). Similarly to the single-layer case, we predict the transmission by assuming decay factor $\alpha=0.8$ within each layer, decay factor $\beta_{1}=0.4$ when the input transmits from one layer to another layer, and $\beta_{2}=1$ when two layers have input and output at the same location, like the output $L_{\rm T}$ for the input $I_{\rm T}$ (Fig. 4d). We find good agreement between the theoretical predictions and the experimental measurements (Fig. 6).
Figure 6: Experimental outputs of the three-layer metamaterial with single inputs. a, Left layer. b, Right layer. Inputs are applied to the middle layer. Solid lines are the theoretical predictions.
Figure 7: 3D printed metamaterials and experimental method. a, Metamaterial for buckling with elastoplastic material (Nylon). b, Metamaterial with tracking markers at the connected joints. c, Stretching stress-strain curve of 3D printed Nylon where we linear fit the elastic modulus and tangent modulus. The white stickers on the dogbone sample are used to track the strain. d, A single-layer metamaterial for local actuation is printed with two materials. Purple ligaments and red diamonds are printed with soft material (Agilus) and rigid material (Vero), respectively. e, Experimental setup of a three-layer metamaterial with local actuation under filming of two cameras. The picture in the top-right corner is the recorded image from the right camera
I.3 Experimental Methods
The metamaterials with floppy modes designed in this work consist of a plurality of triangular unit cells, wherein the rigid bonds and rotating nodes from the theoretical model are replaced with solid diamond blocks of length $\ell$ and width $d$ and ligaments of thickness $t$, respectively. For the metamaterials with buckling floppy modes, we 3D printed them with a single material of $\ell=10.3$ mm, $t=0.4$ mm, and $d=3$ mm (Fig. 7a). For the metamaterials with yield buckling, we used an elastoplastic material, nylon (with a low ratio of $E_{t}/E=0.56\%$ see calibration in Fig. 7c), with the fused deposition modelling (FDM) technology (Ultimaker 2S). For the metamaterials with elastic buckling, we 3D printed them with a rubber-like polymer material (Stratasys, Agilus+Vero). The thicknesses of the metamaterials are set to $D=24.5$ mm for elastoplastic metamaterials and $D=23.2$ mm for elastic metamaterials to prevent out-of-plane buckling. We then marked the ligaments on one side of the metamaterials with white dots for image tracking (Fig. 7b). We used Image-J software to track the position of the white dots of each frame in the video. For the metamaterials with local actuation, we changed the design of ligaments from a rectangular section to a section with inserted tapered rigid parts (Fig. 7d). We then printed such ligaments with a soft material (Agilus) and the diamonds with a rigid material (Vero). Such ligament design can significantly increase resistance to shear and tension while offering only a negligible resistance to bending [39]. For the single-layer metamaterial (Fig. 4b), we set the thickness of the sample to 6 mm. For the three-layer metamaterials with multiple inputs and outputs, we printed each layer separately and assembled them with two rigid frames. The middle layer has a thickness of 6 mm and the left and right layers have a thickness of 3.6 mm. We set a small distance of 0.5 mm between each layer to avoid friction during actuation. We then fixed the sample on the table. We marked the ligaments at the six edges of the hexagon layer with black dots (Fig. 7e top-right corner). We then manually pulled and pushed the rigid bars of the middle layer and recorded the movement of the 12 dots from the left and right layers with two Basler cameras (Fig. 7e).

### v3

@@ I.1 Bounds on Number of Floppy Modes
To maximize the number $F$ of floppy modes, we minimize the number $R$ of rigid chains and maximize the number $L$ of loops. The minimal value of $R$ is zero. We now derive two upper bounds on $L$. The first stems from the fact that each loop surrounds at least one internal corner node, thus $L\leq\frac{N+2-P}{2}$, which leads to $F\leq N_{1}+1$.
@@
For given lattice size and shape, the system is more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with E1E-1 bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in the expression for the number of modes, we obtain the lower bound given in Eq. (1).
@@
For high fractions $N_{1}/N$ of $T_{1}$ blocks, random lattices are not likely to possess loops, thus $L\ll N$, and clearly also $L-R\ll N$. Thus, the typical number of floppy modes is close to the lower bound, as seen in Fig. 1d.
Figure 5: Average number of floppy modes in random lattices. a, Results for different system sizes normalized as $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, so that the lower bounds coincide. b, Normalized number of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
I.2 Number of Floppy Modes in Layered Systems
For a metamaterial made of $K$ coupled layers, each with $N$ triangles and perimeter $P$, the total number of edge nodes is $\left(\frac{3N}{2}+\frac{P}{2}\right)K$, and the total number of bonds between them is $N_{1}+2N_{2}+C$, where $C$ is the number of inter-layer connectors. Note that $N_{1}$ and $N_{2}$ are the total numbers of $T_{1}$ and $T_{2}$ triangles across all layers. Thus, their sum is given by $N_{1}+N_{2}=NK$. Following the same counting arguments presented above for the single-layer case, we now obtain $F=N_{1}-\frac{NK}{2}+\frac{PK}{2}-C+L-R$, where $L$ and $R$ count loops and rigid chains that may now also close by crossing between multiple layers.
I.3 Matrix-Vector Multiplication
For a single input $x$ or $y$, we denote the linear response between the input and outputs as $\begin{pmatrix}u&v\end{pmatrix}^{T}=A\begin{pmatrix}x&y\end{pmatrix}^{T}$ and $\begin{pmatrix}x_{0}&y_{0}\end{pmatrix}^{T}=H\begin{pmatrix}x&y\end{pmatrix}^{T}$, where $x_{0}$ and $y_{0}$ are the displacements generated at the other output node.
To find the matrix $M$ relating all input and output displacements $\vec{U}=\begin{pmatrix}x&y&u&v\end{pmatrix}^{T}$, we consider the reaction forces at the inputs and outputs, $\vec{F}=\begin{pmatrix}f_{x}&f_{y}&f_{u}&f_{v}\end{pmatrix}^{T}$. The compliance matrix $\mathcal{C}$ relates them by $\vec{U}=\mathcal{C}\vec{F}$.
For a single input $x$, $\vec{F}=\begin{pmatrix}f_{x}&0&0&0\end{pmatrix}^{T}$, thus $\vec{U}=\begin{pmatrix}\mathcal{C}_{xx}&\mathcal{C}_{yx}&\mathcal{C}_{ux}&\mathcal{C}_{vx}\end{pmatrix}^{T}f_{x}$. However, in this case we may also write $\vec{U}=\begin{pmatrix}1&H_{yx}&A_{ux}&A_{vx}\end{pmatrix}^{T}x$. Therefore, $\vec{\mathcal{C}}_{x}\equiv\begin{pmatrix}\mathcal{C}_{xx}&\mathcal{C}_{yx}&\mathcal{C}_{ux}&\mathcal{C}_{vx}\end{pmatrix}^{T}=\begin{pmatrix}1&H_{yx}&A_{ux}&A_{vx}\end{pmatrix}^{T}\mathcal{C}_{xx}$. For a single input $y$, we similarly obtain
$\vec{\mathcal{C}}_{y}\equiv\begin{pmatrix}\mathcal{C}_{xy}&\mathcal{C}_{yy}&\mathcal{C}_{uy}&\mathcal{C}_{vy}\end{pmatrix}^{T}=\begin{pmatrix}H_{xy}&1&A_{uy}&A_{vy}\end{pmatrix}^{T}\mathcal{C}_{yy}$.
For two inputs $x$ and $y$, $\vec{F}=\begin{pmatrix}f_{x}&f_{y}&0&0\end{pmatrix}^{T}$, and $\vec{U}=\begin{pmatrix}\vec{\mathcal{C}}_{x}&\vec{\mathcal{C}}_{y}\end{pmatrix}\begin{pmatrix}f_{x}&f_{y}\end{pmatrix}^{T}$. Namely,
$\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}\mathcal{C}_{xx}&\mathcal{C}_{xy}\\ \mathcal{C}_{yx}&\mathcal{C}_{yy}\end{pmatrix}\begin{pmatrix}f_{x}\\ f_{y}\end{pmatrix}$ and $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\mathcal{C}_{ux}&\mathcal{C}_{uy}\\ \mathcal{C}_{vx}&\mathcal{C}_{vy}\end{pmatrix}\begin{pmatrix}f_{x}\\ f_{y}\end{pmatrix}$. We rewrite these two relations as,
$\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\mathcal{C}_{ux}&\mathcal{C}_{uy}\\ \mathcal{C}_{vx}&\mathcal{C}_{vy}\end{pmatrix}\begin{pmatrix}\mathcal{C}_{xx}&\mathcal{C}_{xy}\\ \mathcal{C}_{yx}&\mathcal{C}_{yy}\end{pmatrix}^{-1}\begin{pmatrix}x\\ y\end{pmatrix}=M\begin{pmatrix}x\\ y\end{pmatrix}$. Now, $\begin{pmatrix}\mathcal{C}_{ux}&\mathcal{C}_{uy}\\ \mathcal{C}_{vx}&\mathcal{C}_{vy}\end{pmatrix}\begin{pmatrix}\mathcal{C}_{xx}&\mathcal{C}_{xy}\\ \mathcal{C}_{yx}&\mathcal{C}_{yy}\end{pmatrix}^{-1}=\begin{pmatrix}A_{ux}&A_{uy}\\ A_{vx}&A_{vy}\end{pmatrix}\begin{pmatrix}\mathcal{C}_{xx}\\ \mathcal{C}_{yy}\end{pmatrix}\begin{pmatrix}1&H_{xy}\\ H_{yx}&1\end{pmatrix}^{-1}\begin{pmatrix}\mathcal{C}_{xx}\\ \mathcal{C}_{yy}\end{pmatrix}^{-1}$ . Therefore, $M=AH^{-1}$.
The same method can be generalized to any number of inputs $\vec{X}=\begin{pmatrix}x_{1}&x_{2}&...&x_{n}\end{pmatrix}$ and outputs $\vec{O}=\begin{pmatrix}u_{1}&u_{2}&...&u_{m}\end{pmatrix}$, leading to $\vec{O}=A_{m\times n}H_{n\times n}^{-1}\vec{X}$.
I.4 Matrices Design Space
To theoretically calculate the matrices $A$ and $H$ based on the metamaterial’s design, we assume the displacement decays by a factor $\alpha$ in each bar of the metamaterial. The factor $\alpha$ is determined by the geometry of the metamaterials, which is related to the hinge deformation. Due to frustration, we expect a smaller factor $\alpha$ in the odd loop than in the open chain and even loop. In each triangular unit, if the path from input to output includes one bar, the direction of displacement is conserved, while if there are two bars, it is flipped. Therefore, for an open chain, $A_{ij}$ and $H_{ij}$ are both given by $(-1)^{t_{ij}+b_{ij}+1}\alpha^{b_{ij}}$, where $t_{ij}$ is the number of triangles along the path between input $i$ and output $j$, and $b_{ij}$ is the number of bonds along this path. For hexagons such as those given in Fig. 4, the numbers of triangles are $t_{xy}=2$, $t_{xu}=t_{yv}=3$, $t_{xv}=t_{yu}=4$.
For hexagons with all possible open chains connecting $x$ and $y$, the numbers of bonds may take the values $b_{xy}=6,7,8,9$, $b_{xu},b_{yv}=3,4,5$, and $b_{xv},b_{yu}=4,5,6,7$. For the specific design given in Fig. 4a, this leads to,
$A_{open}=\begin{pmatrix}\alpha^{4}&-\alpha^{6}\\ -\alpha^{4}&\alpha^{4}\end{pmatrix}$, $H_{open}=\begin{pmatrix}1&-\alpha^{8}\\ -\alpha^{8}&1\end{pmatrix}$.
For a metamaterial with a loop, there are two paths between the input and output. If the loop is even, the two paths lead to output displacements in the same direction, and we assume that the magnitude of this displacement is dominated by the shorter of the two paths. For hexagons with even loops, as defined in the Main Text, the possible numbers of bonds along the shorter path are, $b_{xy}=2,3,4$, $b_{xu},b_{yv}=3,4,5$, and $b_{xv},b_{yu}=4,5,6$, and for the design given in Fig. 4b, we have, $A_{even}=\begin{pmatrix}\alpha^{4}&-\alpha^{4}\\ \alpha^{5}&-\alpha^{5}\end{pmatrix}$, $H_{even}=\begin{pmatrix}1&-\alpha^{2}\\ -\alpha^{2}&1\end{pmatrix}$.
In a metamaterial with an odd loop, the two paths lead to output displacements in opposite directions. We assume that these add up such that $A_{ij}$ and $H_{ij}$ are given by the sum of $(-1)^{t_{ij}+b_{ij}+1}\alpha^{b_{ij}}$ with $t_{ij}$ and $b_{ij}$ of the two paths. The design space of a hexagon with an odd loop has, $b_{xy}=(2,7),(2,9),(3,6),(3,8),(4,7)$, $b_{xu},b_{yv}=(3,6),(3,8),(4,7),(5,6)$, and $b_{xv},b_{yu}=(4,5),(4,7),(5,6)$, where pairs indicate the values of $b_{ij}$ along the two paths. Tracking the two paths from each input to each output node for the structure of Fig. 4c, we obtain $A_{odd}=\begin{pmatrix}\alpha^{4}-\alpha^{7}&\alpha^{5}-\alpha^{6}\\ -\alpha^{4}+\alpha^{5}&\alpha^{4}-\alpha^{5}\end{pmatrix}$,
$H_{odd}=\begin{pmatrix}1&\alpha^{3}-\alpha^{8}\\ \alpha^{3}-\alpha^{8}&1\end{pmatrix}$.
See Supplementary Material for realizations of all possible values of $b_{ij}$ listed above.
I.5 Sequential Buckling vs. Simultaneous Buckling
We show that sequential yield buckling with arbitrarily many orders can be generalized in any other elastoplastic materials with periodically arranged floppy modes (Fig. 6a-f), whereas in elastic metamaterials, floppy modes buckle simultaneously, regardless of length differences (Fig. 6g-i).
Figure 6: Sequential actuation of floppy modes. a, Metamaterial with three identical floppy modes printed of plastic Nylon. b, Compression of the metamaterial at different strokes. Color-coded arrows in b, e, and h indicate horizontal displacement. c, Corresponding vertical force and average horizontal displacement, $u_{1}$ of each floppy chain vs. global compressing stroke $u/H$. d, Metamaterial with the same design as panel a, printed of stainless steel. e and f, Corresponding compression at three different strokes and force-displacement curves. g, 3D printed elastic metamaterial with three floppy chains of varying length. h and i, Elastic metamaterial at different stages of compression and force-displacement curves. $B$, $M$, and $T$ in c, f, and i denote the bottom, middle, and top floppy chains. The sign of the slope ($S$) of force vs. compression is detected (dashed lines).